Grade 12th passElectrostatics

Please please please explain this question to me I am in great need for this.A point charge Q is placed at the centre of a hollow uniformly charged sphere of radius R and charge Q.The energy of the system is
Ans-3KQ2/2R

Kaushki
7 Years agoGrade 12th pass

1 Answer

Kapil Khare
7 Years ago
Total energy of the system = Interaction energy of sphere and point charge + Self energy of the Hollow sphere
 
Self energy of the hollow sphere containing the charge Q = KQ2/(2R)
Potential at the centre of sphere due to charge on hollow sphere = KQ/R
\implies Interaction energy between sphere and the point charge = KQ2/R
 
\implies Total  energy of the system = KQ2/(2R) + KQ2/R
\implies T.E. = 3KQ2/(2R)